OpenStudy (anonymous):

this person just posted a question "And ive been tryna work it out ohhhhh" phych seriously tho rlly i even need help tryna figure it out ..........The formula to find the surface area of a sphere is to take the radius of the sphere and square it and multiply that amount by pi and then multiply that by 4. This is written algebraically as S=4πr2, where S is the surface area, r is the radius, and the value of π is 3.14. Rewrite the equation to solve for the radius of the sphere if you know the sphere’s surface area. Then estimate the radius of a sphere given a surface area of 500 meters2.

OpenStudy (anonymous):

and who ever answers this please show me the steps to show me were i went wrong

OpenStudy (dumbcow):

solve for r \[4 \pi r^{2} = S\] \[r^2 = \frac{S}{4\pi}\] \[r = \sqrt{\frac{S}{4\pi}} = \frac{\sqrt{S}}{2 \sqrt{\pi}}\]

OpenStudy (anonymous):

wow i got nothing naer that

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i was using the wrong method

OpenStudy (anonymous):

oops o-0

OpenStudy (dumbcow):

yw what method?? its just algebra

OpenStudy (anonymous):

ik tht method....

OpenStudy (anonymous):

oh i nbwas going to put ty fot thank you but i didnt kno if u knew text language

OpenStudy (anonymous):

in algebra theres diff set ups

OpenStudy (anonymous):

and thats what my brain farts at